OpenStudy (anonymous):

write the equation in standard form: 1/3y= x + 3

OpenStudy (paki):

you know what is standard form...?

OpenStudy (paki):

ok... please have a look here... it will help you a lot.... http://www.algebra-class.com/writing-equations-2.html

OpenStudy (paki):

Standard form is C = aX + bY....

OpenStudy (anonymous):

standard form is Ax+By=C.

OpenStudy (paki):

@shereenkhan yeah same as mine one...

OpenStudy (anonymous):

okay :)

OpenStudy (paki):

no...

OpenStudy (anonymous):

1/3y= x + 3 3x - y = -9

OpenStudy (paki):

standard form is the simplest one.... in general.... @CO_oLBoY

OpenStudy (anonymous):

first multiply the equation by 3 to get rid of the fraction

OpenStudy (anonymous):

@paki i know i corrected that :P

OpenStudy (anonymous):

thats what i did @kelliegirl33 :P

OpenStudy (anonymous):

I see that, however, you didn't explain it....I thought the asker should know how you got the answer

OpenStudy (anonymous):

mhm good :)

OpenStudy (anonymous):

I meant no disrespect

OpenStudy (anonymous):

neither do i

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)

OpenStudy (paki):

@CO_oLBoY be cool :)

OpenStudy (anonymous):

i am cooooooooooool :P :P

OpenStudy (anonymous):

just having fun :)

OpenStudy (anonymous):

every one is smiling here.

OpenStudy (anonymous):

u should too @shereenkhan

OpenStudy (paki):

@shereenkhan yeah, it's because i am here.... :)

OpenStudy (anonymous):

lol @paki

OpenStudy (anonymous):

to smile there must be reason @CO_oLBoY

OpenStudy (anonymous):

its good to smile so keep smiling @shereenkhan

OpenStudy (anonymous):

I am smiling because I am alive...how's that for a reason...lol

OpenStudy (paki):

reason is me :/ @shereenkhan

OpenStudy (paki):

now smile :D

OpenStudy (anonymous):

lol...that just made my day

OpenStudy (anonymous):

paki thats you???

OpenStudy (anonymous):

:P :P

OpenStudy (paki):

@CO_oLBoY no... this is my puppy :/ i am a human :/

OpenStudy (anonymous):

oh good to know i thought may be u r....... :P :P

OpenStudy (anonymous):

they say some dogs have a resemblence to their masters....lol...what about you Paki ?

OpenStudy (anonymous):

@kelliegirl33 u r right that more then enough to smile :)

OpenStudy (anonymous):

i hate dogs :P

OpenStudy (anonymous):

@kelliegirl33 i think this dog has a lot of resemblence to his master XD XD :P

OpenStudy (anonymous):

i don't think so @ coolBoY

OpenStudy (anonymous):

@shereenkhan i hate them when they run after someone XD

OpenStudy (anonymous):

@shereenkhan i know i'm just kidding :P

OpenStudy (anonymous):

How can you hate dogs....there is unconditional love there

OpenStudy (anonymous):

but I guess your right..there is no unconditional love when their chewing your butt off...lol

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

Yes there r nice. but i don't knw i hate them.

OpenStudy (anonymous):

so your a cat person ?

OpenStudy (anonymous):

nope. i am a bird person.

OpenStudy (anonymous):

i have 2 pet cats still i'm not a cat person lol

OpenStudy (anonymous):

I have 2 dogs and a cat....I used to have a canary, but it didn't work out...lol

OpenStudy (anonymous):

my cat Rosco and tweety didn't get along too well

OpenStudy (anonymous):

i have love birds. they are so sweet and cute.

OpenStudy (anonymous):

@shereenkhan this if 4 u

OpenStudy (anonymous):

thank U @CO_oLBoY :)

OpenStudy (anonymous):

thats sweet....well I got to go now...my cereal is getting mushy. Catch you guys later :)

OpenStudy (anonymous):

okay :) it was nice to talk u @kelliegirl33 . u r so amazing person :)

OpenStudy (anonymous):

ttyl @kelliegirl33

OpenStudy (anonymous):

welcome :) @shereenkhan

OpenStudy (anonymous):

:) thanks peeps

OpenStudy (anonymous):

welcome @kelliegirl33 :)

OpenStudy (goformit100):

Warm Welcome To Open Study ......... Sir/ Ma'am @diovannywinster.hermitanio.3

OpenStudy (paki):

@kelliegirl33 sorry, i have no dog :/

OpenStudy (paki):

@CO_oLBoY salay etna free na ho :/

OpenStudy (paki):

@shereenkhan good :/

OpenStudy (anonymous):

@paki acha bhai tum to ghusa hi kr ge

OpenStudy (paki):

@CO_oLBoY am ok....

OpenStudy (anonymous):

. . . but I digress . . .